Solve: 1.3y + 4.4 = 0.8y + 4.1
Found 2 solutions by fractalier, checkley71:
Answer by fractalier(6550)

From 1.3y + 4.4 = 0.8y + 4.1
we subtract 0.8y from both sides and get
0.5y + 4.4 = 4.1
Now subtract 4.4 from both sides and get
0.5y = -0.3
Now divide both sides by 0.5 and get
y = -0.6

Answer by checkley71(8403)

1.3y+4.4=.8y+4.1 or .5y=-.3 y=-3/5
1.3(-3/5)+4.4=.8(-3/5)+4.1 or -3.9/5+4.4=-24/5+4.1 or -.78+4.4=-.48+4.4 or
3.62=3.62